OpenStudy (erinweeks):

A statement Sn about the positive integers is given. Write statements Sk and Sk+1, simplifying Sk+1 completely.

OpenStudy (erinweeks):

Sn: 1 + 4 + 7 + . . . + (3n - 2) = n(3n - 1)/2

OpenStudy (anonymous):

What do you need help with?

OpenStudy (anonymous):

i.e. which part is confusing you?

OpenStudy (erinweeks):

figuring this out?

OpenStudy (erinweeks):

all of it, it confuses me can you do step by step with me?

OpenStudy (anonymous):

No prob.

OpenStudy (anonymous):

Does the question say anything about "Induction"?

OpenStudy (anonymous):

The first part is easy: just look at the "Sn" part and replace the "n"s with "k"s.

OpenStudy (anonymous):

with me?

OpenStudy (erinweeks):

okay with you

OpenStudy (anonymous):

So you're there?

OpenStudy (erinweeks):

yes

OpenStudy (anonymous):

Because you seem to be, um, preoccupied...

OpenStudy (erinweeks):

no im here sorry i was just messaging my teacher quick

OpenStudy (anonymous):

Ok, so you're ready?

OpenStudy (erinweeks):

yes

OpenStudy (anonymous):

You sure

OpenStudy (erinweeks):

yes positive

OpenStudy (anonymous):

Ok. First, explain to me why you're confused. Also, explain to me what you think the question is saying and what it's asking you... Then I'll walk you through step by step.

OpenStudy (erinweeks):

im confused because im not sure exactly how to break it down . & I think its asking me to simplify Sk & Sk+1

OpenStudy (anonymous):

First, notice how the series starts at 1 and then each term that follows is equal the previous term plus 3. 1st term = 1 2nd term = 1+3 = 4 3rd term = 4+3 = 7 4th term = 7+3 = 10 ... Right?

OpenStudy (erinweeks):

okay im following

OpenStudy (anonymous):

Also notice how the "k-th" term in the series: kth-term = 3k -2 where "k" can be any positive whole number. Thus, for the 5th term, we can either calculate it as: 5th-term = 4th-term +3 = 10+3 = 13 Or, we can calculate it as: 5th-term = 3(5) -2 = 13 Either way, we get the same result....

OpenStudy (erinweeks):

okay!

OpenStudy (anonymous):

So you get what I'm saying?

OpenStudy (erinweeks):

yea im on your page

OpenStudy (anonymous):

So that's what the left side of your equation was saying: Take the first "n" numbers of that form and add them all together...

OpenStudy (erinweeks):

okay!

OpenStudy (anonymous):

By the way, in your picture, what's that thing behind you? For a while I thought it was a sock puppet you were holding (weird!). Anyway, I digress...

OpenStudy (erinweeks):

Its my feet !!

OpenStudy (anonymous):

No, the black thing--I swear it looked to me like your shoulder was really your forearm holding up a black hand puppet (Note: I might be crazy).

OpenStudy (erinweeks):

its a coat on a chair lol

OpenStudy (anonymous):

Ok, so anyway, if I asked you to sum up the first 7 numbers of that form, you could easily do it on your calculator, right? You would just add: 1+4+7+10+13+16+19 = 70

OpenStudy (anonymous):

Easy, right?

OpenStudy (anonymous):

But what if I asked you to add up the first 1000 numbers of that form--that would be a real pain in the neck. And that's where the right-side comes in.

OpenStudy (erinweeks):

the first part was easy lol !

OpenStudy (anonymous):

What do you mean?

OpenStudy (erinweeks):

meaning adding the first 7 numbers , but 1000 would be a pain

OpenStudy (anonymous):

So the right side tells us that the sum of the first n numbers of that form is equal to: \[\frac{ n(3n-1) }{ 2 }\] So, when n=7, we get 7(20)/2 = 70 (the same thing we got when we added things up individually) and we can do that (much simpler) calculation for any "n" we want.

OpenStudy (anonymous):

Got it?

OpenStudy (erinweeks):

got it. so thats it?

OpenStudy (anonymous):

What do you mean?

OpenStudy (erinweeks):

was that all i had to do?

OpenStudy (anonymous):

Yeah, the last part of the question is easy: you just replace the "n"s with "k"s and then, for the second part, just replace the n's with "k+1"s and simplify. The question itself is easy, but that's because its trying to slowly introduce something that's more complicated (so it's good that you understand how it all works now).

OpenStudy (anonymous):

Also, something you'll see pretty soon is called "Sigma notation"--it's a way of concisely representing that sum on the lefthand side. It looks like this: \[S _{n}=\sum_{k=1}^{n}(3k-2)\] What it's saying is "start at k=1 and add all the terms from 'k' up to 'n'"

OpenStudy (erinweeks):

so Sk: (3k - 2) = k( 3k -1) /2 or do i plot 7 in for k now and then in sk+1 simplify it all?

OpenStudy (anonymous):

No, because \[S _{k} = 1 + 4 + 7 + 10 + .... + 3k-2\] (i.e. not just "3k-2" but all the terms from 1 to "3k-2", counting by threes) And that sum is equal to \[\ S _{k} = \frac{ k(3k-1) }{ 2 }\]

OpenStudy (erinweeks):

so do i put exactally what you put for sk or do i put what we did individually

OpenStudy (anonymous):

what do you mean?

OpenStudy (erinweeks):

meaning for sk= do i put 1+ 4 + 7 _ 10+... + 3k -2 or do i put 1st term = 1 2nd term = 1+3 = 4 3rd term = 4+3 = 7 4th term = 7+3 = 10

OpenStudy (anonymous):

You write it the first way ( 1+4+7+...+3k-2) The second way was just me showing you where each term came from and what that whole Sn junk was all about...

OpenStudy (erinweeks):

okay dokie

OpenStudy (anonymous):

and remember, you want 1+4+...+3k-2 = k(3k-1)/2 ie. Don't forget the second half!

OpenStudy (anonymous):

and for k+1-th term, it's just 1+4+...+3(k+1)-2, which simplifies to ...+3k+1 etc.

OpenStudy (erinweeks):

Sk+1: 1+4+7+...+3(k+1)-2 = + 3k+1 Thatd be the second half right?

OpenStudy (anonymous):

You mean for S_k+1?

OpenStudy (anonymous):

that would be the lefthand side of the second half

OpenStudy (anonymous):

DOn't forget the righthand side

OpenStudy (anonymous):

1+4+...+3k+1

OpenStudy (erinweeks):

okay

OpenStudy (anonymous):

Then the right hand side being "(k+1)(3(k+1)-1)/2" and then simplify

OpenStudy (anonymous):

Whew! I'm exhausted!! I hope you appreciate this!! :)